OpenStudy (anonymous):

A carpenter has a stick that is 84 centimeters long she cuts off 25% then she cuts the stick into 6 equal pieces

OpenStudy (jdoe0001):

ok... noted

OpenStudy (texaschic101):

25% of 84 = 0.25 * 84 = 21 so the carpenter cuts off 21 cm and is left with (84 - 21) = 63 cm now the carpenter cuts the stick (which is now 63 cm) into 6 equal pieces.. 63/6 = 10 1/2 (or 10.5) cm the 6 pieces are cut into 10 1/2 cm (or 10.5 cm)